Micro Problem Assume a straight-line, downward sloping inverse demand curve, p = 100 - q and a Marginal Costs = 20. 1. What iCan I please get help solving this problem? Thank You
Answer #1

Since allocative efficiency occurs when P=MC.

In monopoly profit-maximizing level, Demand(MB) curve is greater than the MC, therefore a monopoly firm does not achieve allocative efficiency.

To allocative efficiency condition P-MC 10- Q-2 Q=100-20 80 units Substituting into price equation P=10c-80 P = 20 TR= p.Q (0

8 - no units. Substituting a into price equation P =100-40 = $60 Dut - Area of D 1 x 3 x H cose (80-bo) (60-20) 6 | DWL = o.s